velmalovesshaggy145:

negative 2/(negative 2 minus 3)^2 graphing rational functions for ap precalc Currently looking for ordered pair and then asymptotes

velmalovesshaggy145:

Its a simple question really

hannahj:

then you should be able to answer it

velmalovesshaggy145:

@hannahj wrote:
then you should be able to answer it
It's a fraction that does not seem as a possible intercept for the hole of the rational function I have to graph.

hannahj:

ok what grade are you in

velmalovesshaggy145:

12th

hannahj:

im only in 10th

velmalovesshaggy145:

@hannahj wrote:
im only in 10th
Got it

velmalovesshaggy145:

@hannahj wrote:
im only in 10th
I was right.

hannahj:

right about what?

Luigi0210:

Your question doesn't make sense, that's probably why. "negative 2/(negative 2 minus 3)^2" is literally just \(\Large \frac{-2}{-2-3} =\frac{-2}{-5} \) You can't even plot that as you'd need an ordered pair. Think you're missing your variable.

Luigi0210:

Oops, I meant: \[\large \frac{-2}{(-2-3)^2} = \frac{-2}{(-5)^2} =\frac{-2}{25} \] but still the same problem.

Stringcheese272:

idk